OpenStudy (anonymous):

Solve for x: |x − 2| + 10 = 12

OpenStudy (mathstudent55):

1. Subtract 10 from both sides . 2. Then turn it into two simple equations by letting what is inside the abs val signs = the right side (eq 1), and letting what is inside the abs val sign = the opposite of the right side (eq 2) 3. Solve both equations.

OpenStudy (anonymous):

i found 0

OpenStudy (mathstudent55):

That's part of it. Here's the whole thing following the steps I wrote above. 1. |x - 2| + 10 = 12 | x - 2| = 2 2. x - 2 = 2 or x - 2 = -2 x = 4 or x = 0

OpenStudy (anonymous):

oh ok thxx, u deverfve a medal :)

OpenStudy (mathstudent55):

Thanks.

OpenStudy (anonymous):

welc